# Latent space models for grouped multiplex networks

### v1

@@ C.1 Dependency of estimation errors on the cosine similarities of the components
In this section, we present experiments exploring how the latent component estimation error depends on their cosine similarities. Similarly to Section 5, we generate latent components following Algorithm 2 with $n=200$, $M=16$, $K=4$, and $s_{v,u}=s_{w,u}=0.1$ unless one of these two parameters is varied. Here, we only present the results for the Gaussian edge-entry distribution, as the error trends are very similar between the two model types. In Figure 7, we vary $s_{v,w},s_{w,u},s_{w,w},s_{u,u}\in[0.1,0.2,\ldots,0.9]$, where cosine similarities, which were not yet introduced, are defined similarly to (24).
@@
Figure 7: Dependency of ARFE on the cosine similarities between different types of latent components. Unless one of the parameters is varied, the networks are generated according to Algorithm 2 with $M=16$ layers on $n=200$ nodes, $K=4$ balanced groups, and $s_{v,w}=s_{w,u}=0.1$.
As the most important general trend, we can see that an increase in any of the similarities either does not affect or improves the error of $\Theta$. Indeed, a high similarity of latent components makes it harder to separate them, but improves the effective sample size needed to estimate their sum because of the similarity of the summed terms. As an illustrative corner case, when all components are perfectly collinear with each other, it becomes sufficient to fit just one low-rank matrix to get an accurate estimator of $\Theta$, even though separation of the latent components becomes impossible in this case due to violation of the identifiability condition (see Proposition 2.1).
Probably, the most unexpected behavior can be observed in the plot of $s_{w,u}$, showing that individual and group latent component errors improve slightly before reaching $s_{w,u}=0.4$ and then rapidly grow for larger $s_{w,u}$. We believe that the initial error improvement occurs because, by definition, $s_{w,u}$ induces correlation between each group component and individual components not only within but also outside of the corresponding group. Based on the group-wise separation structure of the first stage of Algorithm 1, it is clear that the correlation of a group component with its own individual components makes their separation harder, while the correlation with other components can induce additional information, which implies that the trend change around $s_{w,u}=0.4$ is due to reaching the critical point in the trade-off between the within-group and outside-of-the-group correlations.
Another interesting artifact is the inverted U-shape dependency of the shared and group errors from $s_{v,w}$. This can be explained by the fact that for $s_{v,w}$ close to one, group components add information about the latent space basis of the shared component due to their significant overlap. However, because of the difference in the bases of the group components themselves, the overlap with the shared component does not lead to their inferior estimates. As another corner case, for $s_{v,w}$ close to zero, the group components contain no additional information on the shared one. However, each of them can be easily separated from the shared component by orthogonality. On the other hand, the closer $s_{v,w}$ approaches the middle ($s_{v,w}\approx 0.5$) between these two cases, the more difficult the separation task becomes since both simplifying assumptions are no longer valid.
@@
In the $s_{w,w}$ and $s_{v,w}$ plots, we can observe that the similarity of the group components does not affect the individual error $R$. This is expected since the separation of individual components occurs in the first stage and only involves the sum of the shared and group components, which is not affected by the correlation of the additive terms defining it.
It is also worth mentioning that the errors are robust to the increase in both $s_{w,w}$ and $s_{u,u}$, with the estimation accuracy of $S$ and $Q$ starting to decrease only for $s_{w,w}\geq 0.4$ and $s_{u,u}\geq 0.6$. This observation aligns with the form of the theoretical bounds in Theorem 4.1, suggesting that $s_{w,w}$ and $s_{u,u}$ only play a role in the component errors if they, respectively, dominate ${1\over K}\vee n^{1/2-\tau}$ and ${1\over m_{k}}\vee n^{1/2-\tau}$. While the individual component error bound in Theorem 4.1 does not suggest direct dependency of $R$ on $s_{u,u}$, we can see that $s_{u,u}$ appears in the bound on $r^{(I)}$ derived in Proposition 4.1.
C.2 Method comparison for the Logistic edge-entry model.
@@
Here, we repeat the method comparison experiment in Section 5.3, but this time, under the logistic network-generating model. Figure 8 presents the corresponding results. We omit the points for $n=100$ and $M=8$ due to the instability of the Logistic model fitting for smaller data samples.
@@
Figure 8: Change in ARFE of $\Theta$ (upper row) and $S$ (lower row) with the increase of (left column) the number of nodes $n$ with $M=16$ and (right column) the number of layers $M$ with $n=200$. In all simulations, the number of groups is $K=4$ and the latent dimension is $d=3$. Layers are sampled from the Logistic edge-entry model.
Based on the upper row of Figure 8, we can see that with both $n$ and $M$ varying, GroupMultiNeSS is very close to matching the oracle rates while permanently dominating the MultiNeSS model. Similar to the Gaussian results in Figure 3, we can see that the COSIE model is significantly inferior to both the MultiNeSS and GroupMultiNeSS models.
Looking at the lower row, we can see that contrary to the Gaussian model, GroupMultiNeSS exhibits higher error in estimating the shared component $S$ in the logistic case, with both models significantly less accurate than the oracle estimator. This behavior is related to the fact that latent space models with a non-linear link function, such as the sigmoid, produce adjacency matrices of much higher rank than the original latent space and thus usually suffer from overestimated ranks. In our case, the more latent components are estimated simultaneously, the more freedom the model has to overestimate their latent dimensions and thus overfit during the refitting step.
C.3 Additional real data analysis plots.
In this section, we present additional plots mentioned in the real data experiment described in Section 6. In particular, the upper row of Figure 9 shows the four leading latent dimensions of the shared component, and the lower row contains the first two dimensions of the group components aligned with the same Procrustes rotation as the one used in Figure 4.
Figure 9: First four dimensions of the shared component and first two dimensions of the group components, all fitted with the GroupMultiNeSS (all dimensions in the plots are disassortative).

### v2

@@ C.1 Dependency of estimation errors on the cosine similarities of the components
In this section, we present additional experiments exploring how the latent component estimation error depends on their cosine similarities. Similarly to Section 5, we generate latent components by Algorithm 2 with $n=200$, $M=16$, $K=4$, $s_{v,u}=s_{w,u}=0.1$ and vary one of the cosine similarities $s_{w,w},s_{u,u}$ over the grid $0.1,0.2,\ldots,0.9]$ while keeping the others fixed at 0.1. Figure 7 presents the results for the Gaussian edge distribution; the logistic link results are very similar.
@@
Figure 7: Dependency of ARFE on the cosine similarities between group and individual components. Unless one of the parameters is varied, the networks are generated according to Algorithm 2 with $M=16$ layers on $n=200$ nodes, $K=4$ balanced groups, and $s_{v,u}=s_{w,u}=0.1$.
In general, we see that an increase in any of the similarities has a very small effect on the error of $\Theta$. We conjecture that higher similarity of latent components primarily makes their separation harder but has less influence on the effective sample size for estimating their sum.
@@
In the $s_{w,w}$ plot, we can observe that the similarity of the group components affects the separation of the shared and group components, but does not affect the individual error $R$. This is expected since the separation of individual components occurs in the first stage and only involves the sum of the shared and group components, which is not affected by the correlation of the additive terms defining it.
On the other hand, in the $s_{u,u}$ plot, we can observe that the similarity of the individual components affects the errors of all other components. This is also expected as the separation of the individual components occurs in the first stage and thus propagates the errors into the second stage, where separation of shared and group components takes place. A smaller change in the group error $Q$ compared to the shared $S$ may be explained by the fact that the estimation of each group component depends only on the individual components within its group, while the estimation of the shared component depends on all individual components. Notice that this can also be seen from our theoretical bounds in Theorem 4.1, which states that the error of $S$ depends on the maximal similarity $s_{u,u}$ of all individual components and the error of $Q_{k}$ depends on $s_{u,u}^{(k)}$, the maximal similarity of individual components in group $k$.
C.2 Results for the binary edge model
@@
Here, we repeat the methods comparison experiment in Section 5.3 under the logistic edge model. Figure 8 presents the corresponding results. We omit results for $n=100$ and $M=8$ due to the instability of the Logistic model fitting for small samples.
@@
Figure 8: Change in ARFE of $\Theta$ (top row) and $S$ (bottom row) with the increase of (left column) the number of nodes $n$ with $M=16$ and (right column) the number of layers $M$ with $n=200$. In all simulations, the number of groups is $K=4$ and the latent dimension is $d=3$. Layers are sampled from the Logistic edge-entry model.
When it comes to estimating the overall parameter $\Theta$ (top row of Figure 8), GroupMultiNeSS is again very close to the oracle and outperforms both MultiNeSS and COSIE in all regimes; MultiNeSS also does substantially better than COSIE. However, in the estimation of the shared component $S$ (bottom row) there is a bigger gap between GroupMultiNeSS and the oracle, and in fact MutliNeSS outperform GroupMultiNeSS. This is likely due to the fact that the non-linear link function leads to an adjacency matrix with a much higher estimated rank than that of the original latent space, leading to a larger number of noise eigenvalues inflated during the refitting step. This suggests more careful tuning is needed for the logistic link function model, and possibly a modification of the refitting step; we leave this topic for future work.